Page tree
Skip to end of metadata
Go to start of metadata

Attendance: F. Auchère, U. Schühle, T. Grundy, O Ergincan, S. Sidher, D. Williams.

 

 

  1.     CMS ops

Since the last meeting (4th June) there has been the end of the 3rd Remote-Sensing Checkout Window, and there have been 2 scheduled TCM windows. Since the 29th of June, there have been no Remote-sensing science operations, other than STIX (which is always on).

 

 

  1.     Anomalies (if any)

None since last meeting.


 

  1.     CMS data and interpretation 
  2. a) See IAS CMS data tool

 

 Operations times shown for CMS


LTP03:

  1. STP-151: 2021-06-07T00:00 – 03:00 (DOY 158) around WOL
  2. STP-151: 2021-06-09T16:00 – 19:00 (DOY 160) around SA relubrication
  3. STP-152: 2021-06-13T23:00 – 2021-06-14T-03:30 (DOY 164 – 165) around:
    • WOL
    • HGA pointing
  4. STP-153: 2021-06-21T00:00 – 03:00 (DOY 172) around WOL


LTP04:

  1. STP-154: 2021-06-29T03:20 – 09:20 (DOY 180) around TCM 
  2. STP-154: 2021-07-04T01:00 (DOY 185) – STP-155: 2021-07-05T11:30 (DOY 186) SA rotation, Thermogravimetric Analysis, and WOL 
    • 0100 Switch-on
    • 0130-0300 Test mode
    • 0300-0800 CMS-2 TGA
      • level ended up at was slightly lower than we went in on (75).
    • -> 80degC set-point, 1degC/minute controlled heating and cooling
    • 0800-1200 Test mode
    • 1200-1400 Test mode (SA_ROT from 0 to 30 degrees)
    • 1400-2025 CMS-1 TGA + Peltier
      • CMS had an initial set point that was able to cool, because the set point wasn’t perfect (can’t be, and depends on HIS power status)
      • There shouldn’t be much material collected there, but the level does look higher.
      • -> 80degC set-point, 1degC/minute controlled heating and cooling
    • -> End at -10degC, and stay in Peltier control for ~2 hours
    • -> Note: previous test was at -30degC (November 2020), but the set-point could not be maintained. 
    • Also, the S/C is ~10degC warmer now due to heater changes, so the set-point has been raised by 20degC for this attempt. 
    • 2025-2100 Peltier Off
    • 2100-0000 Test mode (monitoring thermal recovery)
    • 0000-0600 Test mode (WOL)
    • 0600-1130 Test mode (monitoring thermal recovery)
    • 1130 switch-off
    • Sunday 4th July (STP154):
    • Monday 5th July (STP155)
    • A better exercise this time around, given what we learned first time around. The completion of the temperature change went almost exactly where it was expected. So we can run Peltier at  –10ºC without issues.
    • LTP05 should be a little colder
  1. STP-156: 2021-07-12T08:50 – 18:00 (DOY193)


Request to Rad to look at this w.r.t the previous TGAs. To be discussed once analysis is complete.


  1.     Highlights/recommendations
  • CMS operations to be performed weekly, capturing interesting events where possible. 
  • CMS Operator should switch on CMS >30 minutes earlier than the time at which the observations would start, particularly if the observations to catch (for example) a WOL would need to start at midnight. CMS operations in SOOP Kitchen are timed such that they start 30 minutes after assumed switch-on. 
  • Note that From LTP04 onwards (i.e., starting from STP-154), SOC will no longer indicate CMS operations in the EFECS with an RSW_EXT_SPICE. They will still be indicated with CMS “observations”, however. 



 

  1. Future operations:

Already planned

 

 Upcoming planned CMS Ops:

 

  1. STP-154: 2021-07-19T16:00 – 19:00 (DOY 200) around SA relubrication
  2. STP-158: 2021-07-25T22:05 – 2021-07-26T07:15 (DOY 206–207) around pre-VGAM TCM opportunity, and WOL
  3. STP-159: 2021-08-01T19:05 – 2021-08-02T13:25 (DOY 213–214) around pre-VGAM TCM opportunity, and WOL
  4. STP-159: 2021-08-04T22:30 – 2021-08-05T01:30 (DOY 216–217) around SA Rotation
  5. STP-160: 2021-08-04T22:30 – 2021-08-05T01:30 (DOY 216–217) around SA Rotation
  6. STP-161: 2021-08-08T19:25 – 2021-08-09T08:35 (DOY 220–221) around TCM and VGAM (2021-08-09T02:00 – 08:35)



To be planned

 

LTP05 — the planning period from 2021-09-27 until 2021-12-27


The Science Operations Working Group held a kick-off meeting this Monday, with the aim of laying out the orbital context and platform constraints for this 3-month planning period, the last LTP of Cruise Phase. Operations will be coordinated among the payload instruments next week, and the plan should be filled in by the end of next week. 


Tim already has a proposal for the CMS operations in the draft plan we’re currently using in SOOP Kitchen


            https://solarorbiter.esac.esa.int/soopkitchen/#/planning/child-plan/LTP05%20Childplan%20for%20SOWG16%20preparation



Request to run CMS during the EGAM itself if power allows.



  1. CMS meeting communication 


Udo put the question about how to inform people in instrument teams about the status of the CMS measurements.


In the past this hasn’t really been communicated beyond the CMS ops team e-mail list, which includes the Spacecraft Ops Manager (José-Luis Pellon) and the Mission Manager (Luis Sánchez).


Proposal: We could stage the monthly meeting minutes on SOC Public such that instrument teams are able to review the status and comments by the team? This is something we already do for (for example) the Low-Latency Pipeline Engineering telecons.



  1.   AOB

 

            TG: SWA operations have been interrupted quite frequently:

  1. seems heater operations have now been agreed: Summary briefing requested
  2. will SWA ops still be prone to interruptions? Affects heaters and therefore SPICE. What is the outlook for SWA stability? [AI on DW]


 

  1. Next meeting

            Post-VGAM meeting — doodle poll to be sent.

 

            https://esait.webex.com/webappng/sites/esait/meeting/info/3919f6334ad24db09786edaab90c5e2a_20210722T150000Z?launchApp=true

_______________________

 

Action Items

 

AI-12: MS to provide CMS1 & CMS2 FoV Open (DW will ask MS) important for next VGAM

AI-13: David Williams to investigate SWA temperatures 28th Jan Open

AI-16: OE will contact Airbus to see what inputs they need from CMS and SOC to compare with their thermal model

            Contacted ADS: Awaiting an answer when the responsible person is back from holidays.

AI-18: SWA operations resuming from STP 146 TBC — David Williams to check exact date/time  because of effect on CMS-2 temperature.

            PAS started operations at DOY 124 04:45.

AI-25: Rad and Orcun to take a look at the TGA data when possible, and compare with previous TGAs.

  • No labels